# LLMs Can’t Handle Peer Pressure: Crumbling under Multi-Agent Social Interactions

### v2

@@ C.2 Effect on different categories of datasets
Qwen2.5-3B Base 47.60% 47.01% -0.60% 40.24% 42.04% +1.80% 32.13% 36.64% +4.50%
Empowered 57.49% 43.11% -14.37% 53.45% 41.44% -12.01% 63.36% 34.23% -29.13%
Reflected 47.60% 47.60% +0.00% 40.24% 42.04% +1.80% 32.13% 31.53% -0.60%
SFT 58.08% 46.71% -11.38% 43.24% 33.03% -10.21% 27.03% 23.12% -3.90%
GRPO-MAS-DS-DR 54.79% 50.60% -4.19% 57.06% 51.65% -5.41% 55.56% 53.75% -1.80%
GRPO-MAS-NS-OR 65.27% 52.69% -12.57% 63.06% 52.55% -10.51% 64.86% 60.06% -4.80%
GRPO-nonMAS-DS-DR 63.47% 49.70% -13.77% 62.46% 54.35% -8.11% 56.46% 40.24% -16.22%
GRPO-nonMAS-NS-OR 68.56% 51.20% -17.37% 63.36% 52.55% -10.81% 62.76% 60.06% -2.70%
GRPO-MAS-DS-DR-LConf 55.69% 47.31% -8.38% 49.55% 46.25% -3.30% 51.65% 48.95% -2.70%
GRPO-MAS-DS-DR-LCorr 53.89% 47.01% -6.89% 49.85% 45.95% -3.90% 58.56% 55.26% -3.30%
GRPO-MAS-NS-OR-LConf 61.68% 46.41% -15.27% 55.86% 48.05% -7.81% 65.77% 57.36% -8.41%
GRPO-MAS-NS-OR-LCorr 61.98% 45.21% -16.77% 61.56% 47.15% -14.41% 64.86% 57.96% -6.91%
Qwen2.5-7B Base 58.68% 51.20% -7.49% 49.25% 46.85% -2.40% 31.23% 39.04% +7.81%
Empowered 60.78% 50.00% -10.78% 76.58% 50.75% -25.83% 66.97% 44.44% -22.52%
Reflected 58.68% 49.10% -9.58% 49.25% 50.45% +1.20% 31.23% 55.26% +24.02%
SFT 64.97% 43.71% -21.26% 54.65% 39.34% -15.32% 28.83% 36.34% +7.51%
GRPO-MAS-DS-DR 71.26% 66.77% -4.49% 62.46% 61.56% -0.90% 66.67% 57.36% -9.31%
GRPO-MAS-NS-OR 76.65% 74.25% -2.40% 73.57% 70.57% -3.00% 71.17% 73.87% +2.70%
GRPO-nonMAS-DS-DR 71.26% 59.58% -11.68% 63.96% 62.46% -1.50% 62.16% 57.66% -4.50%
GRPO-nonMAS-NS-OR 83.53% 65.27% -18.26% 78.68% 64.86% -13.81% 81.38% 67.57% -13.81%
GRPO-MAS-DS-DR-LConf 65.27% 56.89% -8.38% 51.65% 56.46% +4.80% 54.65% 52.85% -1.80%
GRPO-MAS-DS-DR-LCorr 62.57% 49.10% -13.47% 59.16% 47.75% -11.41% 59.76% 54.95% -4.80%
GRPO-MAS-NS-OR-LConf 72.16% 56.29% -15.87% 75.98% 58.26% -17.72% 76.28% 68.47% -7.81%
GRPO-MAS-NS-OR-LCorr 72.46% 52.40% -20.06% 75.38% 61.26% -14.11% 76.58% 69.67% -6.91%
Qwen2.5-14B Base 62.57% 59.28% -3.29% 62.46% 55.56% -6.91% 40.24% 43.84% +3.60%
Empowered 62.87% 61.38% -1.50% 71.17% 59.16% -12.01% 75.08% 72.97% -2.10%
Reflected 62.57% 59.28% -3.29% 62.46% 54.35% -8.11% 40.24% 49.55% +9.31%
SFT 70.96% 44.91% -26.05% 68.77% 36.04% -32.73% 35.14% 36.34% +1.20%
GRPO-MAS-DS-DR 82.04% 82.63% +0.60% 78.68% 66.67% -12.01% 72.67% 69.07% -3.60%
GRPO-MAS-NS-OR 86.53% 82.34% -4.19% 83.48% 73.57% -9.91% 78.08% 74.77% -3.30%
GRPO-nonMAS-DS-DR 80.24% 72.16% -8.08% 75.38% 65.17% -10.21% 66.37% 66.67% +0.30%
GRPO-nonMAS-NS-OR 86.53% 73.05% -13.47% 86.49% 70.27% -16.22% 80.48% 72.37% -8.11%
GRPO-MAS-DS-DR-LConf 74.85% 59.58% -15.27% 69.07% 59.76% -9.31% 60.96% 60.06% -0.90%
GRPO-MAS-DS-DR-LCorr 75.45% 44.91% -30.54% 66.37% 47.15% -19.22% 62.16% 51.95% -10.21%
GRPO-MAS-NS-OR-LConf 78.44% 61.98% -16.47% 81.98% 66.07% -15.92% 79.28% 66.07% -13.21%
GRPO-MAS-NS-OR-LCorr 78.44% 54.19% -24.25% 77.18% 65.47% -11.71% 75.68% 71.17% -4.50%
Llama3.2-3B Base 49.70% 46.71% -2.99% 30.03% 24.02% -6.01% 22.52% 26.73% +4.20%
Empowered 49.40% 44.61% -4.79% 29.13% 25.53% -3.60% 21.62% 22.52% +0.90%
Reflected 49.70% 35.93% -13.77% 30.03% 29.13% -0.90% 22.52% 21.02% -1.50%
SFT 55.09% 38.62% -16.47% 34.53% 35.44% +0.90% 18.92% 35.14% +16.22%
GRPO-MAS-DS-DR 53.59% 44.01% -9.58% 27.33% 34.83% +7.51% 27.63% 30.93% +3.30%
GRPO-MAS-NS-OR 61.38% 51.80% -9.58% 33.63% 39.64% +6.01% 39.34% 42.64% +3.30%
GRPO-nonMAS-DS-DR 62.87% 44.31% -18.56% 33.03% 30.93% -2.10% 39.34% 32.73% -6.61%
GRPO-nonMAS-NS-OR 60.78% 44.91% -15.87% 43.54% 42.64% -0.90% 47.45% 38.74% -8.71%
GRPO-MAS-DS-DR-LConf 55.99% 42.22% -13.77% 29.13% 34.53% +5.41% 30.93% 28.23% -2.70%
GRPO-MAS-DS-DR-LCorr 54.79% 46.11% -8.68% 26.73% 34.83% +8.11% 28.53% 33.63% +5.11%
GRPO-MAS-NS-OR-LConf 63.77% 52.69% -11.08% 31.53% 38.14% +6.61% 46.25% 40.54% -5.71%
GRPO-MAS-NS-OR-LCorr 63.77% 52.69% -11.08% 31.53% 38.14% +6.61% 46.25% 40.54% -5.71%
Llama3.1-8B Base 55.39% 47.31% -8.08% 32.43% 37.84% +5.41% 32.13% 29.43% -2.70%
Empowered 59.88% 46.71% -13.17% 46.25% 40.54% -5.71% 54.35% 28.83% -25.53%
Reflected 55.39% 40.72% -14.67% 32.43% 33.33% +0.90% 32.13% 22.52% -9.61%
SFT 63.47% 36.53% -26.95% 49.55% 39.34% -10.21% 21.02% 33.93% +12.91%
GRPO-MAS-DS-DR 65.57% 59.28% -6.29% 44.44% 40.84% -3.60% 43.84% 38.14% -5.71%
GRPO-MAS-NS-OR 69.16% 62.87% -6.29% 60.36% 43.54% -16.82% 48.95% 37.24% -11.71%
GRPO-nonMAS-DS-DR 64.37% 47.60% -16.77% 47.15% 38.74% -8.41% 34.53% 33.93% -0.60%
GRPO-nonMAS-NS-OR 74.25% 61.98% -12.28% 58.86% 47.45% -11.41% 48.05% 39.94% -8.11%
GRPO-MAS-DS-DR-LConf 65.27% 50.00% -15.27% 40.54% 35.74% -4.80% 33.33% 37.84% +4.50%
GRPO-MAS-DS-DR-LCorr 69.16% 43.41% -25.75% 44.44% 41.14% -3.30% 46.25% 40.54% -5.71%
GRPO-MAS-NS-OR-LConf 68.86% 54.79% -14.07% 54.05% 38.14% -15.92% 50.45% 42.04% -8.41%
GRPO-MAS-NS-OR-LCorr 71.86% 56.59% -15.27% 49.85% 38.74% -11.11% 48.65% 37.84% -10.81%
@@
Qwen2.5-3B Base 54.33% 46.87% -7.46% 57.66% 67.57% +9.91%
Empowered 61.19% 46.27% -14.93% 56.46% 67.87% +11.41%
Reflected 54.33% 48.36% -5.97% 57.66% 60.66% +3.00%
SFT 48.96% 41.19% -7.76% 45.95% 43.24% -2.70%
GRPO-MAS-DS-DR 61.49% 52.84% -8.66% 54.65% 51.65% -3.00%
GRPO-MAS-NS-OR 65.67% 58.81% -6.87% 57.96% 57.06% -0.90%
GRPO-nonMAS-DS-DR 62.39% 52.54% -9.85% 56.16% 55.86% -0.30%
GRPO-nonMAS-NS-OR 64.78% 53.73% -11.04% 55.26% 46.55% -8.71%
GRPO-MAS-DS-DR-LConf 59.10% 50.75% -8.36% 54.95% 50.15% -4.80%
GRPO-MAS-DS-DR-LCorr 65.97% 47.16% -18.81% 54.35% 41.44% -12.91%
GRPO-MAS-NS-OR-LConf 68.06% 52.24% -15.82% 57.06% 50.15% -6.91%
GRPO-MAS-NS-OR-LCorr 66.57% 45.67% -20.90% 55.86% 43.54% -12.31%
Qwen2.5-7B Base 60.00% 51.04% -8.96% 67.27% 63.36% -3.90%
Empowered 63.58% 57.01% -6.57% 66.97% 65.47% -1.50%
Reflected 60.00% 59.40% -0.60% 67.27% 64.56% -2.70%
SFT 54.93% 41.19% -13.73% 52.55% 46.55% -6.01%
GRPO-MAS-DS-DR 68.06% 63.58% -4.48% 63.66% 58.86% -4.80%
GRPO-MAS-NS-OR 71.94% 68.36% -3.58% 60.36% 53.15% -7.21%
GRPO-nonMAS-DS-DR 68.96% 60.00% -8.96% 59.16% 58.56% -0.60%
GRPO-nonMAS-NS-OR 76.42% 62.69% -13.73% 59.76% 53.15% -6.61%
GRPO-MAS-DS-DR-LConf 65.37% 57.01% -8.36% 64.56% 52.85% -11.71%
GRPO-MAS-DS-DR-LCorr 68.96% 52.54% -16.42% 62.46% 44.44% -18.02%
GRPO-MAS-NS-OR-LConf 74.03% 61.19% -12.84% 66.97% 55.26% -11.71%
GRPO-MAS-NS-OR-LCorr 74.03% 57.61% -16.42% 66.07% 41.74% -24.32%
Qwen2.5-14B Base 67.16% 61.19% -5.97% 73.57% 77.48% +3.90%
Empowered 69.55% 63.88% -5.67% 76.28% 77.78% +1.50%
Reflected 67.16% 66.27% -0.90% 73.57% 72.07% -1.50%
SFT 64.48% 42.39% -22.09% 71.17% 58.26% -12.91%
GRPO-MAS-DS-DR 77.31% 68.96% -8.36% 72.97% 69.37% -3.60%
GRPO-MAS-NS-OR 78.51% 67.16% -11.34% 74.47% 70.87% -3.60%
GRPO-nonMAS-DS-DR 77.61% 65.07% -12.54% 69.37% 66.97% -2.40%
GRPO-nonMAS-NS-OR 78.81% 68.66% -10.15% 71.77% 66.07% -5.71%
GRPO-MAS-DS-DR-LConf 72.24% 63.88% -8.36% 73.27% 72.97% -0.30%
GRPO-MAS-DS-DR-LCorr 71.64% 49.55% -22.09% 69.37% 49.85% -19.52%
GRPO-MAS-NS-OR-LConf 79.40% 61.79% -17.61% 72.07% 59.46% -12.61%
GRPO-MAS-NS-OR-LCorr 80.00% 54.93% -25.07% 63.36% 37.84% -25.53%
Llama3.2-3B Base 48.06% 38.21% -9.85% 52.25% 57.96% +5.71%
Empowered 47.46% 39.70% -7.76% 56.16% 61.56% +5.41%
Reflected 48.06% 34.03% -14.03% 52.25% 37.24% -15.02%
SFT 38.81% 34.33% -4.48% 43.24% 36.04% -7.21%
GRPO-MAS-DS-DR 50.15% 44.18% -5.97% 54.65% 48.35% -6.31%
GRPO-MAS-NS-OR 53.73% 48.36% -5.37% 49.55% 50.75% +1.20%
GRPO-nonMAS-DS-DR 52.84% 40.30% -12.54% 57.06% 49.25% -7.81%
GRPO-nonMAS-NS-OR 57.91% 45.97% -11.94% 55.26% 55.86% +0.60%
GRPO-MAS-DS-DR-LConf 50.15% 38.21% -11.94% 54.95% 54.95% +0.00%
GRPO-MAS-DS-DR-LCorr 53.13% 46.27% -6.87% 55.26% 46.55% -8.71%
GRPO-MAS-NS-OR-LConf 55.82% 45.37% -10.45% 50.75% 50.15% -0.60%
GRPO-MAS-NS-OR-LCorr 55.82% 45.37% -10.45% 50.75% 50.15% -0.60%
Llama3.1-8B Base 60.60% 48.66% -11.94% 58.26% 69.37% +11.11%
Empowered 64.48% 51.04% -13.43% 62.16% 68.77% +6.61%
Reflected 60.60% 48.66% -11.94% 58.26% 52.25% -6.01%
SFT 44.48% 37.01% -7.46% 32.13% 34.53% +2.40%
GRPO-MAS-DS-DR 62.09% 52.84% -9.25% 57.36% 57.96% +0.60%
GRPO-MAS-NS-OR 65.37% 51.94% -13.43% 53.75% 58.26% +4.50%
GRPO-nonMAS-DS-DR 60.00% 49.55% -10.45% 56.76% 61.56% +4.80%
GRPO-nonMAS-NS-OR 65.97% 52.84% -13.13% 43.54% 52.85% +9.31%
GRPO-MAS-DS-DR-LConf 61.19% 49.25% -11.94% 58.86% 60.96% +2.10%
GRPO-MAS-DS-DR-LCorr 63.58% 42.99% -20.60% 56.76% 52.55% -4.20%
GRPO-MAS-NS-OR-LConf 64.48% 53.73% -10.75% 54.35% 48.95% -5.41%
GRPO-MAS-NS-OR-LCorr 64.48% 54.33% -10.15% 57.06% 56.76% -0.30%

### v3

@@ Appendix G Comprehensive Evaluation and Analysis Results
Qwen2.5-3B Base 47.60% 47.01% -1.3% 40.24% 42.04% +4.5% 32.13% 36.64% +14.0%
Empowered 57.49% 43.11% -25.0% 53.45% 41.44% -22.5% 63.36% 34.23% -46.0%
Reflected 47.60% 47.60% +0.0% 40.24% 42.04% +4.5% 32.13% 31.53% -1.9%
SFT 58.08% 46.71% -19.6% 43.24% 33.03% -23.6% 27.03% 23.12% -14.4%
GRPO-MAS-DS-DR 54.79% 50.60% -7.6% 57.06% 51.65% -9.5% 55.56% 53.75% -3.2%
GRPO-MAS-NS-OR 65.27% 52.69% -19.3% 63.06% 52.55% -16.7% 64.86% 60.06% -7.4%
GRPO-nonMAS-DS-DR 63.47% 49.70% -21.7% 62.46% 54.35% -13.0% 56.46% 40.24% -28.7%
GRPO-nonMAS-NS-OR 68.56% 51.20% -25.3% 63.36% 52.55% -17.1% 62.76% 60.06% -4.3%
GRPO-MAS-DS-DR-LConf 55.69% 47.31% -15.0% 49.55% 46.25% -6.7% 51.65% 48.95% -5.2%
GRPO-MAS-DS-DR-LCorr 53.89% 47.01% -12.8% 49.85% 45.95% -7.8% 58.56% 55.26% -5.6%
GRPO-MAS-NS-OR-LConf 61.68% 46.41% -24.8% 55.86% 48.05% -14.0% 65.77% 57.36% -12.8%
GRPO-MAS-NS-OR-LCorr 61.98% 45.21% -27.1% 61.56% 47.15% -23.4% 64.86% 57.96% -10.7%
Qwen2.5-7B Base 58.68% 51.20% -12.8% 49.25% 46.85% -4.9% 31.23% 39.04% +25.0%
Empowered 60.78% 50.00% -17.7% 76.58% 50.75% -33.7% 66.97% 44.44% -33.6%
Reflected 58.68% 49.10% -16.3% 49.25% 50.45% +2.4% 31.23% 55.26% +76.9%
SFT 64.97% 43.71% -32.7% 54.65% 39.34% -28.0% 28.83% 36.34% +26.0%
GRPO-MAS-DS-DR 71.26% 66.77% -6.3% 62.46% 61.56% -1.4% 66.67% 57.36% -14.0%
GRPO-MAS-NS-OR 76.65% 74.25% -3.1% 73.57% 70.57% -4.1% 71.17% 73.87% +3.8%
GRPO-nonMAS-DS-DR 71.26% 59.58% -16.4% 63.96% 62.46% -2.3% 62.16% 57.66% -7.2%
GRPO-nonMAS-NS-OR 83.53% 65.27% -21.9% 78.68% 64.86% -17.6% 81.38% 67.57% -17.0%
GRPO-MAS-DS-DR-LConf 65.27% 56.89% -12.8% 51.65% 56.46% +9.3% 54.65% 52.85% -3.3%
GRPO-MAS-DS-DR-LCorr 62.57% 49.10% -21.5% 59.16% 47.75% -19.3% 59.76% 54.95% -8.0%
GRPO-MAS-NS-OR-LConf 72.16% 56.29% -22.0% 75.98% 58.26% -23.3% 76.28% 68.47% -10.2%
GRPO-MAS-NS-OR-LCorr 72.46% 52.40% -27.7% 75.38% 61.26% -18.7% 76.58% 69.67% -9.0%
Qwen2.5-14B Base 62.57% 59.28% -5.3% 62.46% 55.56% -11.1% 40.24% 43.84% +8.9%
Empowered 62.87% 61.38% -2.4% 71.17% 59.16% -16.9% 75.08% 72.97% -2.8%
Reflected 62.57% 59.28% -5.3% 62.46% 54.35% -13.0% 40.24% 49.55% +23.1%
SFT 70.96% 44.91% -36.7% 68.77% 36.04% $\mathbf{-47.6\%}$ 35.14% 36.34% +3.4%
GRPO-MAS-DS-DR 82.04% 82.63% +0.7% 78.68% 66.67% -15.3% 72.67% 69.07% -5.0%
GRPO-MAS-NS-OR 86.53% 82.34% -4.8% 83.48% 73.57% -11.9% 78.08% 74.77% -4.2%
GRPO-nonMAS-DS-DR 80.24% 72.16% -10.1% 75.38% 65.17% -13.5% 66.37% 66.67% +0.5%
GRPO-nonMAS-NS-OR 86.53% 73.05% -15.6% 86.49% 70.27% -18.8% 80.48% 72.37% -10.1%
GRPO-MAS-DS-DR-LConf 74.85% 59.58% -20.4% 69.07% 59.76% -13.5% 60.96% 60.06% -1.5%
GRPO-MAS-DS-DR-LCorr 75.45% 44.91% -40.5% 66.37% 47.15% -29.0% 62.16% 51.95% -16.4%
GRPO-MAS-NS-OR-LConf 78.44% 61.98% -21.0% 81.98% 66.07% -19.4% 79.28% 66.07% -16.7%
GRPO-MAS-NS-OR-LCorr 78.44% 54.19% -30.9% 77.18% 65.47% -15.2% 75.68% 71.17% -5.9%
Llama3.2-3B Base 49.70% 46.71% -6.0% 30.03% 24.02% -20.0% 22.52% 26.73% +18.7%
Empowered 49.40% 44.61% -9.7% 29.13% 25.53% -12.4% 21.62% 22.52% +4.2%
Reflected 49.70% 35.93% -27.7% 30.03% 29.13% -3.0% 22.52% 21.02% -6.7%
SFT 55.09% 38.62% -29.9% 34.53% 35.44% +2.6% 18.92% 35.14% $\mathbf{+85.7\%}$
GRPO-MAS-DS-DR 53.59% 44.01% -17.9% 27.33% 34.83% +27.5% 27.63% 30.93% +11.9%
GRPO-MAS-NS-OR 61.38% 51.80% -15.6% 33.63% 39.64% +17.9% 39.34% 42.64% +8.4%
GRPO-nonMAS-DS-DR 62.87% 44.31% -29.5% 33.03% 30.93% -6.4% 39.34% 32.73% -16.8%
GRPO-nonMAS-NS-OR 60.78% 44.91% -26.1% 43.54% 42.64% -2.1% 47.45% 38.74% -18.4%
GRPO-MAS-DS-DR-LConf 55.99% 42.22% -24.6% 29.13% 34.53% +18.6% 30.93% 28.23% -8.7%
GRPO-MAS-DS-DR-LCorr 54.79% 46.11% -15.8% 26.73% 34.83% $\mathbf{+30.3\%}$ 28.53% 33.63% +17.9%
GRPO-MAS-NS-OR-LConf 63.77% 52.69% -17.4% 31.53% 38.14% +21.0% 46.25% 40.54% -12.3%
GRPO-MAS-NS-OR-LCorr 63.77% 52.69% -17.4% 31.53% 38.14% +21.0% 46.25% 40.54% -12.3%
Llama3.1-8B Base 55.39% 47.31% -14.6% 32.43% 37.84% +16.7% 32.13% 29.43% -8.4%
Empowered 59.88% 46.71% -22.0% 46.25% 40.54% -12.3% 54.35% 28.83% $\mathbf{-47.0\%}$
Reflected 55.39% 40.72% -26.5% 32.43% 33.33% +2.8% 32.13% 22.52% -29.9%
SFT 63.47% 36.53% $\mathbf{-42.5\%}$ 49.55% 39.34% -20.6% 21.02% 33.93% +61.4%
GRPO-MAS-DS-DR 65.57% 59.28% -9.6% 44.44% 40.84% -8.1% 43.84% 38.14% -13.0%
GRPO-MAS-NS-OR 69.16% 62.87% -9.1% 60.36% 43.54% -27.9% 48.95% 37.24% -23.9%
GRPO-nonMAS-DS-DR 64.37% 47.60% -26.1% 47.15% 38.74% -17.8% 34.53% 33.93% -1.7%
GRPO-nonMAS-NS-OR 74.25% 61.98% -16.5% 58.86% 47.45% -19.4% 48.05% 39.94% -16.9%
GRPO-MAS-DS-DR-LConf 65.27% 50.00% -23.4% 40.54% 35.74% -11.8% 33.33% 37.84% +13.5%
GRPO-MAS-DS-DR-LCorr 69.16% 43.41% -37.2% 44.44% 41.14% -7.4% 46.25% 40.54% -12.3%
GRPO-MAS-NS-OR-LConf 68.86% 54.79% -20.4% 54.05% 38.14% -29.5% 50.45% 42.04% -16.7%
GRPO-MAS-NS-OR-LCorr 71.86% 56.59% -21.2% 49.85% 38.74% -22.3% 48.65% 37.84% -22.2%
Llama3.3-70B Base 64.67% 67.37% +4.2% 57.66% 59.46% +3.1% 35.14% 38.74% +10.2%
Empowered 64.37% 69.49% +8.0% 57.66% 58.56% +1.6% 36.34% 39.94% +9.9%
Reflected 64.67% 66.77% +3.2% 57.66% 55.56% -3.6% 35.14% 31.23% -11.1%
Qwen2.5-32B Base 69.16% 67.96% -1.7% 68.17% 64.86% -4.9% 42.34% 43.84% +3.5%
Empowered 67.07% 63.77% -4.9% 69.97% 63.06% -9.9% 55.26% 46.55% -15.8%
Reflected 69.16% 61.98% -10.4% 68.17% 60.36% -11.5% 42.34% 44.74% +5.7%
Qwen2.5-72B Base 68.26% 66.17% -3.1% 64.56% 64.26% -0.5% 42.04% 43.54% +3.6%
Empowered 68.86% 67.66% -1.7% 64.56% 66.37% +2.8% 42.94% 45.65% +6.3%
Reflected 68.26% 65.57% -3.9% 64.56% 63.36% -1.9% 42.04% 45.95% +9.3%
GPT-OSS-120B Base 89.22% 82.93% -7.0% 98.50% 94.89% -4.6% 98.50% 92.49% -6.1%
Empowered 89.22% 87.72% -1.7% 98.80% 96.40% -2.4% 98.50% 93.09% -5.5%
Reflected 89.22% 88.92% -0.3% 98.50% 95.50% -3.0% 98.50% 91.89% -6.7%
Gemini-2.5-Pro Base 92.81% 80.84% -12.9% 97.60% 91.89% -5.9% 99.10% 95.20% -3.9%
Empowered 90.12% 91.92% +2.0% 97.00% 97.30% +0.3% 99.40% 99.40% +0.0%
Reflected 92.81% 93.11% +0.3% 97.60% 97.90% +0.3% 99.10% 99.70% +0.6%
GPT-5 Base 92.81% 91.32% -1.6% 99.10% 98.80% -0.3% 98.80% 98.50% -0.3%
Empowered 91.92% 92.81% +1.0% 99.10% 98.80% -0.3% 98.80% 99.10% +0.3%
Reflected 92.81% 93.11% +0.3% 99.10% 98.80% -0.3% 98.80% 99.10% +0.3%
@@
Qwen2.5-3B Base 54.33% 46.87% -13.7% 57.66% 67.57% +17.2%
Empowered 61.19% 46.27% -24.4% 56.46% 67.87% +20.2%
Reflected 54.33% 48.36% -11.0% 57.66% 60.66% +5.2%
SFT 48.96% 41.19% -15.8% 45.95% 43.24% -5.9%
GRPO-MAS-DS-DR 61.49% 52.84% -14.1% 54.65% 51.65% -5.5%
GRPO-MAS-NS-OR 65.67% 58.81% -10.5% 57.96% 57.06% -1.6%
GRPO-nonMAS-DS-DR 62.39% 52.54% -15.8% 56.16% 55.86% -0.5%
GRPO-nonMAS-NS-OR 64.78% 53.73% -17.0% 55.26% 46.55% -15.8%
GRPO-MAS-DS-DR-LConf 59.10% 50.75% -14.1% 54.95% 50.15% -8.7%
GRPO-MAS-DS-DR-LCorr 65.97% 47.16% -28.5% 54.35% 41.44% -23.8%
GRPO-MAS-NS-OR-LConf 68.06% 52.24% -23.2% 57.06% 50.15% -12.1%
GRPO-MAS-NS-OR-LCorr 66.57% 45.67% -31.4% 55.86% 43.54% -22.0%
Qwen2.5-7B Base 60.00% 51.04% -14.9% 67.27% 63.36% -5.8%
Empowered 63.58% 57.01% -10.3% 66.97% 65.47% -2.2%
Reflected 60.00% 59.40% -1.0% 67.27% 64.56% -4.0%
SFT 54.93% 41.19% -25.0% 52.55% 46.55% -11.4%
GRPO-MAS-DS-DR 68.06% 63.58% -6.6% 63.66% 58.86% -7.5%
GRPO-MAS-NS-OR 71.94% 68.36% -5.0% 60.36% 53.15% -11.9%
GRPO-nonMAS-DS-DR 68.96% 60.00% -13.0% 59.16% 58.56% -1.0%
GRPO-nonMAS-NS-OR 76.42% 62.69% -18.0% 59.76% 53.15% -11.1%
GRPO-MAS-DS-DR-LConf 65.37% 57.01% -12.8% 64.56% 52.85% -18.1%
GRPO-MAS-DS-DR-LCorr 68.96% 52.54% -23.8% 62.46% 44.44% -28.9%
GRPO-MAS-NS-OR-LConf 74.03% 61.19% -17.3% 66.97% 55.26% -17.5%
GRPO-MAS-NS-OR-LCorr 74.03% 57.61% -22.2% 66.07% 41.74% -36.8%
Qwen2.5-14B Base 67.16% 61.19% -8.9% 73.57% 77.48% +5.3%
Empowered 69.55% 63.88% -8.2% 76.28% 77.78% +2.0%
Reflected 67.16% 66.27% -1.3% 73.57% 72.07% -2.0%
SFT 64.48% 42.39% $\mathbf{-34.3\%}$ 71.17% 58.26% -18.1%
GRPO-MAS-DS-DR 77.31% 68.96% -10.8% 72.97% 69.37% -4.9%
GRPO-MAS-NS-OR 78.51% 67.16% -14.4% 74.47% 70.87% -4.8%
GRPO-nonMAS-DS-DR 77.61% 65.07% -16.2% 69.37% 66.97% -3.5%
GRPO-nonMAS-NS-OR 78.81% 68.66% -12.9% 71.77% 66.07% -8.0%
GRPO-MAS-DS-DR-LConf 72.24% 63.88% -11.6% 73.27% 72.97% -0.4%
GRPO-MAS-DS-DR-LCorr 71.64% 49.55% -30.8% 69.37% 49.85% -28.1%
GRPO-MAS-NS-OR-LConf 79.40% 61.79% -22.2% 72.07% 59.46% -17.5%
GRPO-MAS-NS-OR-LCorr 80.00% 54.93% -31.3% 63.36% 37.84% $\mathbf{-40.3\%}$
Llama3.2-3B Base 48.06% 38.21% -20.5% 52.25% 57.96% +10.9%
Empowered 47.46% 39.70% -16.4% 56.16% 61.56% +9.6%
Reflected 48.06% 34.03% -29.2% 52.25% 37.24% -28.7%
SFT 38.81% 34.33% -11.5% 43.24% 36.04% -16.7%
GRPO-MAS-DS-DR 50.15% 44.18% -11.9% 54.65% 48.35% -11.5%
GRPO-MAS-NS-OR 53.73% 48.36% -10.0% 49.55% 50.75% +2.4%
GRPO-nonMAS-DS-DR 52.84% 40.30% -23.7% 57.06% 49.25% -13.7%
GRPO-nonMAS-NS-OR 57.91% 45.97% -20.6% 55.26% 55.86% +1.1%
GRPO-MAS-DS-DR-LConf 50.15% 38.21% -23.8% 54.95% 54.95% +0.0%
GRPO-MAS-DS-DR-LCorr 53.13% 46.27% -12.9% 55.26% 46.55% -15.8%
GRPO-MAS-NS-OR-LConf 55.82% 45.37% -18.7% 50.75% 50.15% -1.2%
GRPO-MAS-NS-OR-LCorr 55.82% 45.37% -18.7% 50.75% 50.15% -1.2%
Llama3.1-8B Base 60.60% 48.66% -19.7% 58.26% 69.37% +19.1%
Empowered 64.48% 51.04% -20.8% 62.16% 68.77% +10.6%
Reflected 60.60% 48.66% -19.7% 58.26% 52.25% -10.3%
SFT 44.48% 37.01% -16.8% 32.13% 34.53% +7.5%
GRPO-MAS-DS-DR 62.09% 52.84% -14.9% 57.36% 57.96% +1.0%
GRPO-MAS-NS-OR 65.37% 51.94% -20.5% 53.75% 58.26% +8.4%
GRPO-nonMAS-DS-DR 60.00% 49.55% -17.4% 56.76% 61.56% +8.5%
GRPO-nonMAS-NS-OR 65.97% 52.84% -19.9% 43.54% 52.85% $\mathbf{+21.4\%}$
GRPO-MAS-DS-DR-LConf 61.19% 49.25% -19.5% 58.86% 60.96% +3.6%
GRPO-MAS-DS-DR-LCorr 63.58% 42.99% -32.4% 56.76% 52.55% -7.4%
GRPO-MAS-NS-OR-LConf 64.48% 53.73% -16.7% 54.35% 48.95% -10.0%
GRPO-MAS-NS-OR-LCorr 64.48% 54.33% -15.7% 57.06% 56.76% -0.5%
Llama3.3-70B Base 67.76% 66.87% -1.3% 73.57% 74.77% +1.6%
Empowered 68.66% 68.66% +0.0% 77.18% 75.08% -2.7%
Reflected 67.76% 64.78% -4.4% 73.57% 75.08% +2.1%
Qwen2.5-32B Base 73.13% 69.25% -5.3% 78.68% 79.58% +1.1%
Empowered 74.93% 69.25% -7.6% 81.08% 81.98% +1.1%
Reflected 73.13% 69.25% -5.3% 78.68% 77.48% -1.5%
Qwen2.5-72B Base 72.24% 71.34% -1.2% 81.38% 83.48% +2.6%
Empowered 72.84% 73.73% $\mathbf{+1.2\%}$ 82.58% 84.98% +2.9%
Reflected 72.24% 71.64% -0.8% 81.38% 78.08% -4.1%
GPT-OSS-120B Base 88.66% 84.78% -4.4% 79.28% 77.78% -1.9%
Empowered 89.85% 86.27% -4.0% 81.68% 81.98% +0.4%
Reflected 88.66% 85.97% -3.0% 79.28% 84.98% +7.2%
Gemini-2.5-Pro Base 92.54% 83.28% -10.0% 89.79% 85.29% -5.0%
Empowered 91.94% 91.04% -1.0% 90.69% 92.49% +2.0%
Reflected 92.54% 91.34% -1.3% 89.79% 92.79% +3.3%
GPT-5 Base 92.84% 92.84% +0.0% 81.68% 87.99% +7.7%
Empowered 91.34% 91.64% +0.3% 84.38% 90.09% +6.8%
Reflected 92.84% 92.24% -0.6% 81.68% 89.49% +9.6%